Find the length of arc of the curve 𝑦 = π‘₯ 3⁄2 /√3 from (0, 0) to (4, √8 / 3 )


1
Expert's answer
2022-06-07T00:24:00-0400

If "x=4, y=8/\sqrt{3}"


"f'(x)=(\dfrac{x^{3/2}}{\sqrt{3}})'=\dfrac{\sqrt{3}}{2}\sqrt{x}"


"L=\displaystyle\int_{0}^{4}\sqrt{1+(f'(x))^2}dx"

"=\displaystyle\int_{0}^{4}\sqrt{1+(\dfrac{\sqrt{3}}{2}\sqrt{x})^2}dx"

"=\dfrac{1}{2}\displaystyle\int_{0}^{4}\sqrt{4+3x}dx"

"=\dfrac{1}{2}[\dfrac{1}{3}(\dfrac{2}{3}(4+3x)^{3/2})]\begin{matrix}
 4\\
 0
\end{matrix}"

"=\dfrac{1}{9}(64-8)=\dfrac{56}{9}"